OpenStudy (calculusxy):

Happy Pi Day!!! Share your moments with math or even pi! What did you learn new today? It's all about math!!!

OpenStudy (calculusxy):

\[\pi\] rocks!!!

jigglypuff314 (jigglypuff314):

314 is in my name ;) and I found this post :P http://openstudy.com/study#/updates/53991178e4b068926b38b231

OpenStudy (alexandervonhumboldt2):

wow

OpenStudy (calculusxy):

At 9:26:53 it will be epic!!! I am not sure if it already happened in your time zones yet, but since I am in New York now, I am waiting for it!

Nnesha (nnesha):

\(\huge\color{green}{{\rm i~<3}\rm~ chocolate~pi}\)

jigglypuff314 (jigglypuff314):

if you're in new york then it would've already past technically :P 9:26:53 in the morning that is xD if you're thinking PM tsk tsk cuz that is technically 21:26:53 not 9:26:53 :(

OpenStudy (calculusxy):

oh I didn't think of it that way...

OpenStudy (calculusxy):

But I was busy during that time, so I couldn't really "celebrate it." That's why I will just consider it as PM.

OpenStudy (anonymous):

I remember in 2nd grade, my teacher put on a vhs about pi. It talked about the ratio between pi and the diameter of a circle, how they found it and it's uses. I also remember thinking that it was boring and who really cares about pi. Now pi and what can be learned from it is slowly taking over my life. Happy pi day!

Nnesha (nnesha):

\[\huge\rm \sqrt{-64} \times\color{green} \pi\] answer : \(\huge\color{green}{{\rm i~ate~pie}}\)

Nnesha (nnesha):

 o^_^o